OpenStudy (anonymous):

The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line? The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line? @Mathematics

OpenStudy (amistre64):

\[\frac{x}{-5}+\frac{y}{-2}=1\] thatll work

OpenStudy (saifoo.khan):

you have 2 points, (-5,0) and (0,-2) use this to find slope.

OpenStudy (anonymous):

these are my options that i came up with y = –x – 5 y = x + 2 y = x – 2 y = –x – 2

OpenStudy (amistre64):

\[\frac{x}{-5}+\frac{y}{-2}=1\] \[\frac{x}{10/-2}+\frac{y}{10/-5}=1\] \[\frac{-2x}{10}+\frac{-5y}{10}=1\] \[\frac{-2x}{10}+\frac{-5y}{10}=\frac{10}{10}\] \[-2x-5y=10\] or the more proper \[2x+5y=-10\] maybe

OpenStudy (amistre64):

those are your options?

OpenStudy (amistre64):

id go last one as a hunch

OpenStudy (anonymous):

yes i dont know how to make fractions

OpenStudy (amistre64):

\[\frac{x}{-5}+\frac{y}{-2}=1\] \[\frac{x(-2)}{-5}+\frac{y(-2)}{-2}=1(-2)\] \[\frac{x(-2)}{-5}+y=-2\] \[\frac{-2x(-5)}{-5}+(-5)y=-2(-5)\] \[-2x-5y=10\] \[2x+5y=-10\] \[5y=-2x-10\] \[y=\frac{-2}{5}x-\frac{10}{5}\] \[y=\frac{-2}{5}x-2\] this is what i tend to get ....

OpenStudy (anonymous):

yeah that was one of my answers but i dont know how to do fractions

OpenStudy (amistre64):

(-2/5) is a good way

OpenStudy (amistre64):

y = (-2/5)x - 2

OpenStudy (anonymous):

ok then thanks